OpenStudy (anonymous):

An object is launched at 19.6 m/s from a height of 58.8 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t2 +19.6t + 58.8. What is the object's maximum height?

OpenStudy (phi):

find the vertex of the parabola (or use calculus)

OpenStudy (anonymous):

man its been long since ive done that...would you mind giving some clues or somthn?

OpenStudy (phi):

is this calculus?

OpenStudy (anonymous):

supose to be algebra

OpenStudy (phi):

the vertex occurs at x= -b/(2a) in y= ax^2 +bx +c

OpenStudy (phi):

match the letters to the numbers in your equation

OpenStudy (phi):

once you find x (as a number), use it to find the height by replacing x with the number in the equation

OpenStudy (phi):

I mean t in this case, rather than x

OpenStudy (anonymous):

so 58.8=-4.9(19.6)^2+19.6(19.6) + 58.8 ?

OpenStudy (phi):

h(t) = -4.9t2 +19.6t + 58.8 this is a parabola (I shot an arrow into the air, it follows a parabola) the peak occurs at t= -b/(2*a) where a and b are the coefficients of the generic parabola f(t)= a*t^2 +b*t +c match up these equations h(t) = -4.9t2 +19.6t + 58.8 f(t) = a*t^2 +b*t + c to find a and b. then use them in t= -b/(2*a) once you find t as a number, use it in h(t) = -4.9t2 +19.6t + 58.8 to find the height at that t

OpenStudy (anonymous):

thanks @phi